The mass of a block of metal reads as 0.56-kg when weighed in air. Once the block is completely submerged in water the same spring scale reads 0.21-kg. What is the density of this metal? Please report your answer in kg/m3.
Answer #1

Buoyant force = 0.56 - 0.21

Fb = 0.35 N

AND

Fb = m*g*density of fluid/density of object

density of water = 1000 kg/m^3

So,

Fb = 0.51*d(water)/d(metal)

0.35 = 0.51*1000/d(metal)

d(metal) = 1457.142 kg/m^3

Density of the metal is 1457.142 kg/m^3